• Suppose you are being interviewed for your dream job. The pay is tricky and there's also a catch: the job is only for 5 months.
  • Your future boss tells you that the first week's pay will be $10 (not per hour, the whole week is $10!)
  • Concerned about your reaction, your future boss adds that your week 4 salary will be $100 (again, not per hour).
  • Knowing how valuable you would be to the company and before you storm out of the office the boss mentions that the pay increases exponentially every week. 
  • Now... would you take this job?
  1. Find the model that describes this salary situation using the variable "t" as the number of weeks.
  2. Week #1 corresponds to "t=0". Week #2 corresponds to "t=1", etc... How much will you get paid in Week 3?
  3. For simplicity, assume that a month has exactly 4 weeks. How much will you get paid during the last week.




1
Expert's answer
2021-11-24T06:16:05-0500

1.

let x be salary per week, then:

"x(t)=10" for "0\le t\le3"

"x(t)=100" for "t=4"

"x(t)=100(1+r)^{t-4}" for "5\le t\le20"

where r is growth rate


2.

"x(3)=\$10"


3.

"x(20)=100(1+r)^{16}"